OpenStudy (anonymous):

Ten is the geometric mean of 4 and what number? 2 40 25 8

Parth (parthkohli):

So basically: \(\Rightarrow 10 = \sqrt{4x} \) \(\Rightarrow 10 = 2\sqrt x\) \(\Rightarrow 5 = \sqrt x \)

Parth (parthkohli):

Square both sides and you get \(x = 5^2\)

OpenStudy (anonymous):

\[GM = \sqrt{a \times b}\] \[4 \times b = 10^2\] Find b from here..

OpenStudy (anonymous):

\[10=\sqrt{4*25}\]

OpenStudy (anonymous):

so its 25?

Parth (parthkohli):

Yes, yes it is.

OpenStudy (anonymous):

sweet :)

Parth (parthkohli):

:)

OpenStudy (anonymous):

thanks

Parth (parthkohli):

You're welcome.

Parth (parthkohli):

*SHAKA*

OpenStudy (anonymous):

lol